OpenStudy (anonymous):

I need help really fast! Please help! How do you graph the following logarithmic function (graphing technology)?

OpenStudy (anonymous):

\[\log_{3} (x)-1\] I just need the function that you type in

OpenStudy (anonymous):

@Hero

hero (hero):

That IS what you type in. You already have what to type in. What calculator are you using?

OpenStudy (anonymous):

geogebra... That's what they told us to use

hero (hero):

log[3]{x} - 1 I think

OpenStudy (anonymous):

is it a straight line? That's what I got

hero (hero):

No, it wouldn't be a straight line

OpenStudy (nincompoop):

all logs or roots have curve lines. you can test it yourself graph the simplest of their forms y = x^2 y = x^3 y = x^5 y = x^(1/2) …. and so on

OpenStudy (psymon):

For the most part, the log graphs are going to be very similar with only minor differences with the base number change. Before any shift occurs, the point (1,0) always exists in a log graph and the y-axis is an asymptote. This is a general log graph:  Then from there, you just apply any shifts your graph my have. If the base is higher, like log base 100 or something, the curve will just be more like this:  Otherwise no huge differences. A graphing utility can just help you get more precision and certain points if needed, but not that must of a necessity for a basic graph ofit. Your function is log base 3 shifted down 1 :3